OpenStudy (anonymous):

a. Is it possible for AB = 10, BC = 2x + 5 and AC = 5x, given that point B is the midpoint? b. Why or why not?

OpenStudy (anonymous):

sure, why not?

OpenStudy (anonymous):

what.....

OpenStudy (anonymous):

Mind giving a bit more context to your question?

OpenStudy (anonymous):

thats the question

OpenStudy (anonymous):

Then no one can help you. You should ask your teacher to tell you what she is talking about.

OpenStudy (anonymous):

or he.

hartnn (hartnn):

if B is the mid-point AB=BC 10=2x+5 can u find x from here ?

OpenStudy (anonymous):

lol, my bad. didn't read the question properly.

OpenStudy (anonymous):

5=2x ?

OpenStudy (anonymous):

therefore x = ?

OpenStudy (anonymous):

2 and 5 dont go in together

OpenStudy (anonymous):



hartnn (hartnn):

2x=5 x=5/2 ok?

OpenStudy (anonymous):

ohhhhh....

OpenStudy (anonymous):

so b will be the midpoint right.

hartnn (hartnn):

wait, we need to check that. if B is midpoint. AC=2AB 5x=2(10) can u find x from here ?

OpenStudy (anonymous):

x=5

hartnn (hartnn):

sure ?

OpenStudy (anonymous):

so now what ?

hartnn (hartnn):

its x=4 ! check.

OpenStudy (anonymous):

what... am confused

hartnn (hartnn):

5x=2(10) 5x=20 x=4 ok?

OpenStudy (anonymous):

now is point the midpoint

hartnn (hartnn):

since 2 values of x DO NOT MATCH.(4 and 5/2) B cannot be the mid-point.

hartnn (hartnn):

did u get that^ ? @lvasquev462